You are working as a Junior Engineer for a small motor racing team. You have been given a proposed mathematical model to calculate the velocity of a car accelerating from rest in a straight line. The equation is: 𝑣(𝑡) = 𝐴 (1 − 𝑒 − 𝑡 𝑡𝑚𝑎𝑥𝑠𝑝𝑒𝑒𝑑)


Ascari A10 5.0 V8 - [2006] t=2.8 tm(400)= 10.36 tmaxspeed= 7.8


derive an equation a(t) for the instantaneous position of the car as a function of time? Identify the acceleration of the car at t = 0s and asymptote of this function as t → ∞?


1
Expert's answer
2022-06-07T12:02:56-0400
"v(t) = A (1-e ^ {- t\/t_{maxspeed}})"


Using the information that "t_{maxspeed}=7.8\ s" we have


"v(t) = A (1-e ^ {- t\/7.8})"

Using the information that "t (0-28 m\/s)" is "2.8\ s" we have



"v(2.8) = A (1-e ^ {- 2.8\/7.8})=28""A=\dfrac{28}{1-e ^ {- 2.8/7.8}}m/s""A=92.836\ m/s"




"v_{max}=v(t_{maxspeed})=A (1-e ^ {- t_{maxspeed}\/t_{maxspeed}})""v_{max}=A(1-e^{-1})""v_{max}=92.836\ m/s\cdot(1-e^{-1})""v_{max}=58.673\ m/s"




"a(t)=\dfrac{dv}{dt}=\dfrac{A}{t_{maxspeed}} (e ^ {- t/t_{maxspeed}})""=\dfrac{92.836\ m/s}{7.8\ s} (e ^ {- t/7.8}), \ 0\ s\le t\le7.8\ s"

The function "a(t)" decreases for "0\ s\le t\le7.8\ s."



"a_{max}=a(0)=\dfrac{92.836\ m/s}{7.8\ s}"




"a_{max}=11.902\ m/s^2"



If "t\to\infin,a\to0\ m/s^2"